# Unifying the Three Algebraic Approaches to the CSP via Minimal Taylor Algebras

### v5

@@ 9.5 Examples from Section 5
Proof.
The algebra is minimal Taylor because $m$ generates a minimal clone (see [35]). The set $C=\{0,1\}$ is an absorbing subuniverse of $\mathbf{A}$ as witnessed by the 4-ary operation
@@
The set $C$ is not a center of $\mathbf{A}$ since for any potentially witnessing relation $R\leq_{sd}\mathbf{A}\times\mathbf{B}$ the subuniverse $D=2+R\leq\mathbf{B}$ satisfies $m(D,B,B)\subseteq D$ (as $m(2,1,0)=2$) and $m(B,D,D)\subseteq D$ (as $m(0,2,2)=2$), so $m(x,y,y)$ witnesses that $D$ is a 2-absorbing subuniverse of $\mathbf{B}$. ∎
@@
Note that the cyclic permutation (0 1 2 3)0123(0\;1\;2\;3)( 0 1 2 3 ) is an automorphism of $\mathbf{A}$, so the same is true for any reduct of $\mathbf{A}$ (equivalently, the graph of the permutation (0 1 2 3)0123(0\;1\;2\;3)( 0 1 2 3 ) is a binary relation on $\mathbf{A}$ which is preserved by $g$). One can easily verify that the binary relation $\mathbf{S}=\operatorname{Sg}_{\mathbf{A}^{2}}\{(0,1),(1,0)\}$ is equal to $\{(x,y)\mid x=y\pm 1\pmod{4}\}$, and that there is a congruence $\theta$ on $\mathbf{S}$ with congruence classes $\{(x,y)\mid x=y+1\pmod{4}\}$ and $\{(x,y)\mid x=y-1\pmod{4}\}$. Additionally, the quotient $\mathbf{S}/\theta$ is isomorphic to the $2$-element minority algebra.
By the above, the operation $f$ must be compatible with the automorphism (0 1 2 3)0123(0\;1\;2\;3)( 0 1 2 3 ), the binary relation $\mathbf{S}$, and the congruences $\alpha$ and $\theta$. Additionally, $f$ must act as the minority operation on $\{0,2\}$ and on $\mathbf{S}/\theta$, and $f$ must act as the majority operation on $\mathbf{A}^{\prime}/\alpha$.
The above constraints on $f$ allow us to determine every value of $f$ in terms of the value of $f(0,0,1)$. To see this, note that since $f$ acts as the minority operation on $\mathbf{S}/\theta$, we can determine the value of $f(a\pm 1,b\pm 1,c\pm 1)$ from the value of $f(a,b,c)$. For instance, we have $f(1,1,0)=f(0,0,1)-1\pmod{4}$ and $f(0,2,1)=f(1,1,0)-1\pmod{4}=f(0,0,1)-2\pmod{4}$. This, together with the fact that $f$ is cyclic, allows us to determine the value of $f(a,b,c)$ for every triple $a,b,c$ not all of the same parity from the value of $f(0,0,1)$. If $a,b,c$ all have the same parity, then $f$ acts as the minority operation on $\{a,b,c\}$. Since $f$ acts as majority on $\mathbf{A}^{\prime}/\alpha$, we have $f(0,0,1)\in\{0,2\}$, so there are only two possibilities for $f$. If $f(0,0,1)=0$, then $f$ is equal to $g$. Otherwise, if $f(0,0,1)=2$, then we have $g(x,y,z)=f(f(x,x,f(x,y,z)),f(y,y,f(x,y,z)),f(z,z,f(x,y,z)))$, so $f$ and $g$ generate the same clone. ∎
@@
Proof.
The algebra $\mathbf{S}=\operatorname{Sg}_{\mathbf{A}^{2}}\{(a,b),(b,a)\}$ has a congruence $\psi$ such that $\mathbf{S}/\psi$ is isomorphic to $(\mathbb{Z}/4,x-y+z)$. Explicitly, the congruence classes of $\psi$ are $\{(a,b),(c,d)\}$, $\{(b,a),(d,c)\}$, $\{(a,d),(c,b)\}$,$\{(b,c),(d,a)\}$, and in this order they correspond to the elements $0,1,2,3$ of $\mathbb{Z}/4$. Abusing notation, we will identify the congruence classes of $\psi$ with the elements of $\mathbb{Z}/4$ in the remainder of the proof.
To prove that this example is minimal Taylor, first note that every pair of elements forms an abelian edge, so the same must be true in any Taylor reduct. Thus by Theorem 6.5, any Taylor reduct $\mathbf{A}^{\prime}$ of $\mathbf{A}$ must have a Mal’cev term $q(x,y,z)$. The restriction of $q$ to any two-element affine subalgebra or quotient must be the minority operation. This fixes the restriction of $q$ to the sets $\{a,c\}$ and $\{b,d\}$, as well as the restriction of $q$ to the quotient $\mathbf{A}/\theta$, and the fact that $\mathbb{Z}/4$ has only one Mal’cev term forces $q$ to act on $\mathbf{S}/\psi=\mathbb{Z}/4$ as $x-y+z$.
@@ 10 Proofs for Section 6: Omitting types
Proof.
Combining the results of [5, 27] and [22], $\mathbf{A}$ has bounded width if and only if no subalgebra of $\mathbf{A}$ has a nontrivial abelian quotient. The proof of Theorem 2.8 in [50] shows that this is also equivalent to having wnu term operations of every arity $n\geq 3$. Thus, (ii), (iii), and (iv) are equivalent. To show that (ii) implies (i), it suffices to observe that if $\mathbf{A}$ contains an abelian edge $(a,b)$ and congruence $\theta$ witnesses that, then the algebra $\operatorname{Sg}_{\mathbf{A}}(a,b)/\theta$ is abelian. Finally, to show that (i) implies (ii), assume that there is a subalgebra $\mathbf{B}$ of $\mathbf{A}$ and a congruence $\theta$ on $\mathbf{B}$ such that $\mathbf{B}/\theta$ is abelian. Then for any $a,b$ from different equivalence classes of $\theta$, the pair $(a,b)$ is an abelian edge. ∎
@@
Proof.
First, to show that (i) implies (ii) let $B\trianglelefteq_{2}\mathbf{A}^{\prime}\leq\mathbf{A}$. If $B$ is a nonempty proper subset of $A^{\prime}$, then Proposition 9.7 gives us a minimal semilattice edge in $\mathbf{A}^{\prime}$ and thus in $\mathbf{A}$, a contradiction.
By Theorem 2.12. of [12] the existence of a 3-edge term operation is equivalent to a so-called 3-cube operation and then (ii) implies (iii) follows from Theorem 4.5 in [46] (see also [48]) which provides a 3-cube operation whenever $\operatorname{Clo}(\mathbf{A})$ is generated by a single ternary operation (which is the case by Theorem 5.24) and no subalgebra of $\mathbf{A}$ has a nontrivial projective subuniverse (which is the case by Proposition 8.2).
From [12] it follows that (iii) implies (iv) – it is proved there that the existence of an edge (or cube) term operation is equivalent to having few subpowers.
Finally, we show that (iv) implies (i). If $\mathbf{A}$ contains a semilattice edge, then by Theorem 5.13(a) it also has a subalgebra term equivalent to a 2-element semilattice. It is known from [12, 43] that a semilattice does not have few subpowers. ∎
@@
Proof.
To see that (i) implies (ii) recall that 3-absorbing subuniverses and centers are the same by Theorem 5.10, that 3-absorbing subuniverses are stable under abelian and semilattice edges by Theorem 5.21, and that 2-absorbing subuniverses are exactly those stable under all the edges by Theorem 5.19. By item (4) in Proposition 5.9, (ii) implies (ii’). The fact that every majority edge defines two disjoint 3-absorbing subuniverses shows that (ii’) implies (i). ∎
@@
If $\mathbf{A}$ is a minimal Taylor algebra without majority edges and $\mathbb{Z}/{2}$-edges, then $\mathbf{A}$ has a commutative binary term operation.
Proof.
By Lemma 4.4 in [4] it is enough to show that for any $a,b\in\mathbf{A}$ there is a term operation $f$ such that $f(a,b)=f(b,a)$, or in other words that the binary symmetric subpower $\operatorname{Sg}_{\mathbf{A}^{2}}((a,b),(b,a))$ of $\mathbf{A}$ contains a tuple of the form $(c,c)$. Observe that, conversely, any symmetric nonempty binary subpower of an algebra with a commutative binary term operation intersects the diagonal.
Let $\mathbf{A}$ be a minimal counterexample to the proposition with respect to $|A|$ and choose $a,b\in A$ such that $S=\operatorname{Sg}\{(a,b),(b,a)\}$ does not intersect the diagonal. Clearly $\operatorname{Sg}(a,b)=A$ by minimality. If $\mathbf{A}$ has a proper congruence $\theta$, then there is some $c\in\mathbf{A}$ such that $S\cap(c/\theta)^{2}\neq\emptyset$ (since $\mathbf{A}/\theta$ is not a counterexample), but then $S\cap(c/\theta)^{2}\leq c/\theta^{2}$ is a binary nonempty symmetric subpower of $c/\theta$ that avoids the diagonal, a contradiction to the minimality of $\mathbf{A}$.
Suppose now that $\mathbf{A}$ has a proper 3-absorbing subuniverse $B$. If $B\cap(B+S)\neq\emptyset$, then there exist $d,e$ such that $(d,e)\in S\cap B^{2}$ (take any $e\in B\cap(B+S)$ and then $d\in B$ such that $(d,e)\in S$), so since $d,e$ generate a proper subalgebra of $\mathbf{A}$ we see by minimality that $\operatorname{Sg}((d,e),(e,d))\leq\mathbf{S}$ contains a diagonal element, a contradiction. On the other hand, if $B\cap(B+S)=\emptyset$, then by item (2) in Proposition 5.12 $E=B\cup(B+S)$ must be a subuniverse of $\mathbf{A}$ with congruence $\theta$ corresponding to the partition $B,B+S$, such that $\mathbf{E}/\theta$ is a two element majority algebra. Any pair in different equivalence classes is then a majority edge, a contradiction.
It remains to deal with the case that $\mathbf{A}$ is simple and $\mathbf{A}$ has no nontrivial 3-absorbing subuniverse. It then follows from Proposition 9.11 that $\mathbf{A}$ is abelian, so $\mathbf{A}$ is term equivalent to an affine Mal’cev algebra of a group isomorphic to $\mathbb{Z}/{p}$. By the absence of $\mathbb{Z}/{2}$-edges, $p$ is odd. Moreover, Corollary 4.5 implies that $S$ is the graph of a bijection $A\to A$ – the graph of an automorphism of $\mathbf{A}$. Since $S$ is generated by $(a,b)$ and $(b,a)$, the automorphism has order $2$. But $|A|=p$ is odd, so $\alpha(c)=c$ for some $c$, therefore $S$ contains $(c,c)$, a contradiction. ∎
@@
That (iii’) implies (i) follows from the fact that the two element majority algebra and the two element affine Mal’cev algebra both have no nontrivial binary operations. ∎
@@
Proof.
To prove that (i) implies (ii) assume for contradiction that $B$ is a nontrivial absorbing subuniverse of $\mathbf{A}^{\prime}\leq\mathbf{A}$. By Corollary 4.13 there exist $b\in B$ and $a\in A^{\prime}\setminus B$ such that $(a,b)$ or $(b,a)$ is a minimal edge. Since $\mathbf{A}^{\prime}$ is $\mathrm{sm}$-free, we get that $(b,a)$ (and $(a,b)$) is necessarily a minimal abelian edge; let $\theta$ be a witnessing congruence of $\operatorname{Sg}(b,a)$. By Theorem 5.21, the set $B$ is stable under abelian edges, therefore each $\theta$-class intersects $B$, in particular there exists $a^{\prime}$ in $B\cap a/\theta$. But then $\operatorname{Sg}(b,a^{\prime})\subseteq B$ is strictly contained in $\operatorname{Sg}(b,a)$, contradicting the minimality of the edge.
Condition (ii) is the property HAF from [8]. By Theorem 1.4 from the same paper $\mathbf{A}$ has a Mal’cev term operation, so (iii) holds.
Finally, if $\mathbf{A}$ has a semilattice or a majority edge $(a,b)$ witnessed by a congruence $\theta$, then the algebra $\operatorname{Sg}_{\mathbf{A}}(a,b)/\theta$ (and thus $\mathbf{A}$) cannot have a Mal’cev term operation by Theorem 5.13. Therefore (iii) implies (i). ∎
@@
Since both a two-element semilattice and an affine Mal’cev algebra do not have a near unanimity term operation, by Theorem 5.13 (iii) implies (i). ∎
@@ 11 Conclusion
Third, both CSP dichotomy proofs [28, 60] require and develop more advanced Commutator Theory [39, 47] concepts and results, while in this paper we have merely used some fundamental facts about the basic concept, the abelian algebra. Is it possible to develop our theory in this direction as well, potentially providing sufficient tools for the dichotomy result? Also, is there a natural concept that would replace thin edges in Bulatov’s approach?
@@
Finally, there is yet another, older, and highly developed theory of finite algebras, the tame congruence theory started in [42]. What are the connections to the theory initiated in this paper?

### v6

@@ 9.5 Examples from Section 5
Proof 9.49.
The algebra is minimal Taylor because $m$ generates a minimal clone (see [CsakanyMinimalClones]). The set $C=\{0,1\}$ is an absorbing subuniverse of as witnessed by the 4-ary operation
@@
The set $C$ is not a center of since for any potentially witnessing relation $R\leq_{sd}$ the subuniverse $D=2+R\leq$ satisfies $m(D,B,B)\subseteq D$ (as $m(2,1,0)=2$) and $m(B,D,D)\subseteq D$ (as $m(0,2,2)=2$), so $m(x,y,y)$ witnesses that $D$ is a 2-absorbing subuniverse of .
@@
Note that the cyclic permutation (0 1 2 3)0123(0\;1\;2\;3)( 0 1 2 3 ) is an automorphism of $\mathbf{A}$, so the same is true for any reduct of $\mathbf{A}$ (equivalently, the graph of the permutation (0 1 2 3)0123(0\;1\;2\;3)( 0 1 2 3 ) is a binary relation on $\mathbf{A}$ which is preserved by $g$). One can easily verify that the binary relation is equal to $\{(x,y)\mid x=y\pm 1\pmod{4}\}$, and that there is a congruence $\theta$ on with congruence classes $\{(x,y)\mid x=y+1\pmod{4}\}$ and $\{(x,y)\mid x=y-1\pmod{4}\}$. Additionally, the quotient is isomorphic to the $2$-element minority algebra.
By the above, the operation $f$ must be compatible with the automorphism (0 1 2 3)0123(0\;1\;2\;3)( 0 1 2 3 ), the binary relation , and the congruences $\alpha$ and $\theta$. Additionally, $f$ must act as the minority operation on $\{0,2\}$ and on , and $f$ must act as the majority operation on $\mathbf{A}^{\prime}/\alpha$.
The above constraints on $f$ allow us to determine every value of $f$ in terms of the value of $f(0,0,1)$. To see this, note that since $f$ acts as the minority operation on , we can determine the value of $f(a\pm 1,b\pm 1,c\pm 1)$ from the value of $f(a,b,c)$. For instance, we have $f(1,1,0)=f(0,0,1)-1\pmod{4}$ and $f(0,2,1)=f(1,1,0)-1\pmod{4}=f(0,0,1)-2\pmod{4}$. This, together with the fact that $f$ is cyclic, allows us to determine the value of $f(a,b,c)$ for every triple $a,b,c$ not all of the same parity from the value of $f(0,0,1)$. If $a,b,c$ all have the same parity, then $f$ acts as the minority operation on $\{a,b,c\}$. Since $f$ acts as majority on $\mathbf{A}^{\prime}/\alpha$, we have $f(0,0,1)\in\{0,2\}$, so there are only two possibilities for $f$. If $f(0,0,1)=0$, then $f$ is equal to $g$. Otherwise, if $f(0,0,1)=2$, then we have $g(x,y,z)=f(f(x,x,f(x,y,z)),f(y,y,f(x,y,z)),f(z,z,f(x,y,z)))$, so $f$ and $g$ generate the same clone.
@@
Proof 9.51.
The algebra has a congruence $\psi$ such that is isomorphic to $(\mathbb{Z}/4,x-y+z)$. Explicitly, the congruence classes of $\psi$ are $\{(a,b),(c,d)\}$, $\{(b,a),(d,c)\}$, $\{(a,d),(c,b)\}$,$\{(b,c),(d,a)\}$, and in this order they correspond to the elements $0,1,2,3$ of $\mathbb{Z}/4$. Abusing notation, we will identify the congruence classes of $\psi$ with the elements of $\mathbb{Z}/4$ in the remainder of the proof.
To prove that this example is minimal Taylor, first note that every pair of elements forms an abelian edge, so the same must be true in any Taylor reduct. Thus by Theorem 6.5, any Taylor reduct $\mathbf{A}^{\prime}$ of $\mathbf{A}$ must have a Mal’cev term $q(x,y,z)$. The restriction of $q$ to any two-element affine subalgebra or quotient must be the minority operation. This fixes the restriction of $q$ to the sets $\{a,c\}$ and $\{b,d\}$, as well as the restriction of $q$ to the quotient $\mathbf{A}/\theta$, and the fact that $\mathbb{Z}/4$ has only one Mal’cev term forces $q$ to act on as $x-y+z$.
@@ 10 Proofs for Section 6: Omitting types
Proof 10.1.
Combining the results of [Barto14:local, Bulatov16:restricted] and [Bulatov04:graph], $\mathbf{A}$ has bounded width if and only if no subalgebra of has a nontrivial abelian quotient. The proof of Theorem 2.8 in [Kozik15:maltsev] shows that this is also equivalent to having wnu term operations of every arity $n\geq 3$. Thus, (ii), (iii), and (iv) are equivalent. To show that (ii) implies (i), it suffices to observe that if $\mathbf{A}$ contains an abelian edge $(a,b)$ and congruence $\theta$ witnesses that, then the algebra $\operatorname{Sg}_{\mathbf{A}}(a,b)/\theta$ is abelian. Finally, to show that (i) implies (ii), assume that there is a subalgebra $\mathbf{B}$ of $\mathbf{A}$ and a congruence $\theta$ on $\mathbf{B}$ such that $\mathbf{B}/\theta$ is abelian. Then for any $a,b$ from different equivalence classes of $\theta$, the pair $(a,b)$ is an abelian edge.
@@
Proof 10.2.
First, to show that (i) implies (ii) let $B\trianglelefteq_{2}\mathbf{A}^{\prime}\leq$. If $B$ is a nonempty proper subset of $A^{\prime}$, then Proposition 9.22 gives us a minimal semilattice edge in and thus in , a contradiction.
By Theorem 2.12. of [Berman10:varieties] the existence of a 3-edge term operation is equivalent to a so-called 3-cube operation and then (ii) implies (iii) follows from Theorem 4.5 in [kazda21:cube-terms] (see also [kearnes17:cube-term-blockers]) which provides a 3-cube operation whenever $\operatorname{Clo}(\mathbf{A})$ is generated by a single ternary operation (which is the case by Theorem 5.24) and no subalgebra of $\mathbf{A}$ has a nontrivial projective subuniverse (which is the case by Proposition 8.2).
From [Berman10:varieties] it follows that (iii) implies (iv) – it is proved there that the existence of an edge (or cube) term operation is equivalent to having few subpowers.
Finally, we show that (iv) implies (i). If $\mathbf{A}$ contains a semilattice edge, then by Theorem 5.13(a) it also has a subalgebra term equivalent to a 2-element semilattice. It is known from [Berman10:varieties, Idziak10:few] that a semilattice does not have few subpowers.
@@
Proof 10.3.
To see that (i) implies (ii) recall that 3-absorbing subuniverses and centers are the same by Theorem 5.10, that 3-absorbing subuniverses are stable under abelian and semilattice edges by Theorem 5.21, and that 2-absorbing subuniverses are exactly those stable under all the edges by Theorem 5.19. By item (4) in Proposition 5.9, (ii) implies (ii’). The fact that every majority edge defines two disjoint 3-absorbing subuniverses shows that (ii’) implies (i).
@@
If is a minimal Taylor algebra without majority edges and $\mathbb{Z}/{2}$-edges, then has a commutative binary term operation.
Proof 10.5.
By Lemma 4.4 in [Barto12:cyclic] it is enough to show that for any $a,b\in\mathbf{A}$ there is a term operation $f$ such that $f(a,b)=f(b,a)$, or in other words that the binary symmetric subpower $\operatorname{Sg}((a,b),(b,a))$ of contains a tuple of the form $(c,c)$. Observe that, conversely, any symmetric nonempty binary subpower of an algebra with a commutative binary term operation intersects the diagonal.
Let be a minimal counterexample to the proposition with respect to $|A|$ and choose $a,b\in A$ such that $S=\operatorname{Sg}\{(a,b),(b,a)\}$ does not intersect the diagonal. Clearly $\operatorname{Sg}(a,b)=A$ by minimality. If $\mathbf{A}$ has a proper congruence $\theta$, then there is some $c\in\mathbf{A}$ such that $S\cap(c/\theta)^{2}\neq\emptyset$ (since is not a counterexample), but then $S\cap(c/\theta)^{2}\leq c/\theta^{2}$ is a binary nonempty symmetric subpower of $c/\theta$ that avoids the diagonal, a contradiction to the minimality of .
Suppose now that $\mathbf{A}$ has a proper 3-absorbing subuniverse $B$. If $B\cap(B+S)\neq\emptyset$, then there exist $d,e$ such that $(d,e)\in S\cap B^{2}$ (take any $e\in B\cap(B+S)$ and then $d\in B$ such that $(d,e)\in S$), so since $d,e$ generate a proper subalgebra of $\mathbf{A}$ we see by minimality that $\operatorname{Sg}((d,e),(e,d))\leq$ contains a diagonal element, a contradiction. On the other hand, if $B\cap(B+S)=\emptyset$, then by item (2) in Proposition 5.12 $E=B\cup(B+S)$ must be a subuniverse of $\mathbf{A}$ with congruence $\theta$ corresponding to the partition $B,B+S$, such that is a two element majority algebra. Any pair in different equivalence classes is then a majority edge, a contradiction.
It remains to deal with the case that is simple and has no nontrivial 3-absorbing subuniverse. It then follows from Proposition 9.30 that is abelian, so is term equivalent to an affine Mal’cev algebra of a group isomorphic to $\mathbb{Z}/{p}$. By the absence of $\mathbb{Z}/{2}$-edges, $p$ is odd. Moreover, Corollary 4.5 implies that $S$ is the graph of a bijection $A\to A$ – the graph of an automorphism of . Since $S$ is generated by $(a,b)$ and $(b,a)$, the automorphism has order $2$. But $|A|=p$ is odd, so $\alpha(c)=c$ for some $c$, therefore $S$ contains $(c,c)$, a contradiction.
@@
That (iii’) implies (i) follows from the fact that the two element majority algebra and the two element affine Mal’cev algebra both have no nontrivial binary operations.
@@
Proof 10.7.
To prove that (i) implies (ii) assume for contradiction that $B$ is a nontrivial absorbing subuniverse of . By Corollary 4.13 there exist $b\in B$ and $a\in A^{\prime}\setminus B$ such that $(a,b)$ or $(b,a)$ is a minimal edge. Since is $\mathrm{sm}$-free, we get that $(b,a)$ (and $(a,b)$) is necessarily a minimal abelian edge; let $\theta$ be a witnessing congruence of $\operatorname{Sg}(b,a)$. By Theorem 5.21, the set $B$ is stable under abelian edges, therefore each $\theta$-class intersects $B$, in particular there exists $a^{\prime}$ in $B\cap a/\theta$. But then $\operatorname{Sg}(b,a^{\prime})\subseteq B$ is strictly contained in $\operatorname{Sg}(b,a)$, contradicting the minimality of the edge.
Condition (ii) is the property HAF from [Barto15:malcev]. By Theorem 1.4 from the same paper $\mathbf{A}$ has a Mal’cev term operation, so (iii) holds.
Finally, if $\mathbf{A}$ has a semilattice or a majority edge $(a,b)$ witnessed by a congruence $\theta$, then the algebra $\operatorname{Sg}_{\mathbf{A}}(a,b)/\theta$ (and thus ) cannot have a Mal’cev term operation by Theorem 5.13. Therefore (iii) implies (i).
@@
Since both a two-element semilattice and an affine Mal’cev algebra do not have a near unanimity term operation, by Theorem 5.13 (iii) implies (i).
@@ 11 Conclusion
Third, both CSP dichotomy proofs [Bulatov17:dichotomy, Zhuk20:dichotomy] require and develop more advanced Commutator Theory [freese:CMcommutator, kearnes:shape] concepts and results, while in this paper we have merely used some fundamental facts about the basic concept, the abelian algebra. Is it possible to develop our theory in this direction as well, potentially providing sufficient tools for the dichotomy result? Also, is there a natural concept that would replace thin edges in Bulatov’s approach?
@@
Finally, there is yet another, older, and highly developed theory of finite algebras, the tame congruence theory started in [hobby88:tct]. What are the connections to the theory initiated in this paper?